OpenStudy (anonymous):

Contestants in a game show are each given three pairs of numbers (for example, 2-3, 5-6, and 6-1). A person in the audience is randomly selected to roll two dice, one white and one red, where white symbolizes the first number in the pair and red the second number. What is the probability that each player has of winning?

OpenStudy (anonymous):

The chance that the first number is that of a player is 1 in 6 For the second number it is also 1 in 6 So the chance for each player of winning is \[\frac{ 1 }{ 6 } \times \frac{ 1 }{ 6 } = \frac{ 1 }{ 36 } \]